# Bilevel ZOFO: Bridging Parameter-Efficient and Zeroth-Order Techniques for Efficient LLM Fine-Tuning and Meta-Training

### v2

@@ A.2 Fine-tuning LLMs for Multitask and Few-Shot Learning
Multi-task learning (MTL) enables a model to handle multiple tasks simultaneously, fostering knowledge transfer between tasks and improving overall efficiency [34, 54]. Typical meta-tuning approaches employ First-Order methods to train autoregressive LLMs on a multitask dataset for various tasks [63, 34, 13]. Zhong et al. [63] apply meta-training to tasks such as hate speech detection, question categorization, topic classification, and sentiment classification. Guo et al. [13] adopt the method from Min et al. [34] for generating stylistic text. While Min et al. [34] focus on enhancing the in-context learning ability of the meta-trained model for multitask learning, Zhong et al. [63] focus on improving zero-shot performance. This approach is particularly valuable in low-resource settings, where collecting large labeled datasets can be costly, as is often the case with medical data. In such environments, few-shot learning—where a model is fine-tuned on a high-resource dataset to quickly adapt to new tasks with minimal data—becomes essential [55]. To address the challenges of multi-task and few-shot learning in natural language processing, several meta fine-tuning methods have been proposed [17, 62, 55, 1]. However, traditional meta fine-tuning approaches, such as MetaICL [34], still require full-model first-order gradient calculations, which become computationally expensive with large language models (LLMs) containing billions of parameters. During training, Min et al. [34] sample a task from the dataset for each iteration to perform in-context learning. In contrast to Zhong et al. [63] and Min et al. [34], our approach uses a bilevel structure: the full LLM is fine-tuned at the upper level, while parameter-efficient fine-tuning (PEFT) models are tuned at the lower level. At test time, we freeze the meta-tuned base model and fine-tune only the PEFT model using a few-shot setup, which is both more cost-effective and efficient. Crucially, Min et al. [34] fine tune the full model with first order methods, while we employ a ZO method in meta-tuning the base model at the upper level. Our approach allows us to bypass the need for backpropagation in the meta-model, significantly reducing computational costs.
@@ A.3 First-order Methods for Bilevel Optimization
Solving a bilevel optimization problem is challenging because the function value in the upper-level objective depends on the optimizer of the lower-level problem. This makes it difficult to compute the gradient of the upper-level objective, also known as the hypergradient. Classical methods require calculating Hessian-vector multiplications to approximate the hypergradient [8, 9, 7, 23, 41, 12, 2, 30]. However, when fine-tuning large language models, this process becomes extremely expensive due to the high computational and memory demands.
Recently, new frameworks for bilevel optimization have been introduced [31, 44, 27, 21, 26, 18, 32]. These methods bypass the need for second-order information by reformulating the bilevel problem as a constrained optimization problem. The constraint is penalized, allowing the problem to be tackled as a minimax problem using only first-order information. These methods significantly reduce computational costs by eliminating the need for second-order information. Nevertheless, when fine tuning LLMs, back propagation for calculating the gradient of an LLM is still too expensive.
Liu et al. [27] and Lu and Mei [31] explore the convergence of their proposed methods to the original bilevel problem, while other approaches only demonstrate convergence to the penalized problem. In this paper, we adapt the method from Lu and Mei [31] to approximate part of the upper-level parameters using a ZO approximation in order to address the challenge posed by the large number of training parameters in large language models. We also provide convergence guarantees for this adapted zeroth-order-first-order method.
@@ Appendix B Theoretical guarantees of Bilevel ZOFO
The following is the low effective rank assumption from [33]. This assumption avoids dimension $d$ in the total complexity. Following [33], we assume here that ${\mathbf{z}}^{k}$ in (7) is sampled from shpere in $\mathbb{R}^{d}$ with radius $\sqrt{d}$ for ease of illustration.
@@
This result matches the complexity in previous ZO minimax algorithm in [53] but solves our bilevel optimization problem (2) and does not depend on the dimensionality $d$ thanks to the efficient rank assumption B.5, providing efficiency guarantee for our algorithm.
@@ C.1 Proofs
(Lemma 1.2.3, Theorem 2.1.8 and Theorem 2.1.10 in [35])
@@
The following is a rephrase of theorem 2 in [31].
@@
Therefore, by Theorem 2 in [31] we have $\mathbb{E}[\|\nabla F({\bm{\theta}},{\mathbf{p}}^{*}({\bm{\theta}}))\|]\leq O(\epsilon)$ and Lemma B.4 is proven. ∎
@@
(Lemma C.2 in [58]) Let $\mathbf{u}$} be uniformly sampled from the Euclidean sphere $\sqrt{d}\mathbb{{\mathbf{s}}}^{d-1}$ and $\mathbf{v}$} be uniformly sampled from the Euclidean ball $\sqrt{d}\mathbb{B}^{d}=\{{\mathbf{x}}\in\mathbb{R}^{d}\mid\|{\mathbf{x}}\|\leq\sqrt{d}\}$. For any function $f({\mathbf{x}}):\mathbb{R}^{d}\to\mathbb{R}$ and $\alpha>0$, we define its zeroth-order gradient estimator as:
@@
The following lemma is from Theorem 1 in [33].
@@ D.1 Single-Task experiments
Following MeZO [33], we evaluate our approach on a range of classification and multiple-choice tasks: BoolQ [3], CB [52], CB [52], COPA [42], ReCoRD: [59],RTE [51], SST2 [51], WiC [37], WinoGrande [43]. In this setting, training and testing are conducted on the same task.
Hyperparameter Search Given resource limitations, we focus on sweeping only the learning rate as the key hyperparameter. For MeZO and first-order PEFT experiments, we explore learning rates from the set $\{1e-2,1e-3,1e-4,1e-5,1e-6\}$. For Bilevel-ZOFO, we sweep both the upper-level and lower-level learning rates: $\text{lr}_{\text{upper}}\in\{1e-4,1e-5,1e-6\}$ and $\text{lr}_{\text{lower}}\in\{1e-2,1e-3,1e-4,1e-5\}$. We perform all experiments in tables 7 and 8 using three random seeds and report the average and standard deviation. We also set $\epsilon=1e-3$, following MeZO [33].
@@ D.1.1 Training
All experiments use the Adam optimizer [20],including baselines and both lower-level and upper-level optimizers. No weight decay was applied, and the models were trained with a constant learning rate schedule. Batch size is set to $16$ for all experiments. We load all models in bfloat16. We find the best performing model based on validation loss and report test results from that checkpoint. We report the test accuracy or F1-score based on the test dataset being imbalanced or not.
@@ D.2 More Results for single task fine tuning
Table 7 presents the detailed test metrics when applying bilevel-ZOFO and baselines to fine-tune OPT-1.3B [60] on a downstream task.
@@
Table 8 demonstrates the results for fine-tuning Llama2-7b [50] on various classification and open-ended generation tasks.
@@ D.2.1 Compare with the two-stage pipeline
The observed performance drop after the second stage is indeed counter-intuitive at first glance. However, it is a limitation of MeZO as it approximates gradients. While further fine-tuning intuitively should improve performance, the inherent noise in gradient approximation can lead to suboptimal updates. This observation is consistent with the fact that MeZO typically requires a significant number of iterations to converge. This is a key contribution of our work: Our approach addresses MeZO’s challenges, such as sensitivity to hard prompts and long convergence times, while outperforming both MeZO and PEFT and maintaining similar memory efficiency. The intuition behind why our method is effective in enhancing both MeZO’s full-model tuning and PEFT is in the nested bilevel structure. This structure encodes more information (as reflected in the training method) from the prompt tuning stage than only treating it as a first stage, thereby providing better guidance for MeZO. In contrast, our bilevel method effectively addresses the issues of MeZO and demonstrates improved performance over both MeZO and the PEFT baseline, even with the same number of ZO iterations. This phenomenon that a bilevel-method is better than a two-staged pipeline is also observed in the later work on diffusion models [47].
The training loss curves for both stages of a two-stage approach and our bilevel framework are provided in Figure 6. When running MeZO in the second stage, the training loss exhibits oscillations and does not show improvement within 500–1000 iterations. This behavior is consistent with findings in the original MeZO [33] paper, which notes that MeZO typically requires much longer to converge—on the order of 100k iterations. The oscillatory behavior observed within the shorter training duration is not surprising due to gradient approximation errors.
@@ D.3.1 Meta-Tasks
Following the methodology of Min et al. [34], we evaluate the performance of bilevel-ZOFO as a fast and efficient meta-learning algorithm. We perform experiments using four of the distinct meta-learning settings outlined in MetaICL [34]: classification-to-classification, non-classification-to-classification, QA-to-QA, and non-QA-to-QA. Each of these meta-learning tasks includes a set of training sub-tasks and a different set of test sub-tasks. The sub-tasks are sourced from CROSSFIT [55] and UNIFIEDQA [19], comprising a total of 142 unique sub-tasks. These sub-tasks cover a variety of problems, including text classification, question answering, and natural language understanding, all in English. Table 10 shows the number of tasks in each training and testing meta-learning setting and the total number of examples in each training task.
@@
See Tables 14 and 15 of MetaICL [34] for a list of all sub-tasks.
@@ D.3.2 Baselines
We use GPT2-Large [40] as the base model for these experiments.We compare our method against several baseline approaches:
  • MetaICL [34]: A method for meta-learning with in-context learning. MetaICL tunes all the parameters of the base model using the first-order method. In both training and testing, the model is given $k$ demonstration examples, $(a_{1},b_{1}),\dots,(a_{k},b_{k})$}, where $b_{i}$ represents either classification labels or possible answers in question-answering tasks, along with one test example $(a,b)$. The input is formed by concatenating the demonstration examples $a_{1},b_{1},\dots,a_{k},b_{k},a$. The model then computes the conditional probability of each label, and the label with the highest probability is selected as the prediction.
@@
We sample 768 examples from each training sub-task. We use these samples to train MetaICL in their original setting for 30,000 steps. This includes learning rate of 1e51e-5, batch size of $1$ on $8$ GPUs, 8-bit Adam optimizer and fp16 half precision. See MetaICL[34] for full details. To train our method, we split the training dataset of each sub-task to two subsets, 256 samples as the development dataset for upper-level updates and 512 samples for lower-level training. For each outer iteration of our method, we randomly sample a subset of 5 training tasks. We perform 10 lower-level updates between each pair of upper-level updates. To keep bilevel-ZOFO as lightweight as possible, unlike MetaICL, we do not include demonstration examples in the inputs. Since bilevel-ZOFO uses significantly less memory and has much faster updates compared to MetaICL, theoretically we are able to train it for many more iterations within the same total training duration as MetaICL. However, due to resource constraints, we only train bilevel-ZOFO for 50,000 iterations. Similar to  [33], we did not observe a plateau in performance for bilevel-ZOFO, indicating that further training can yield additional improvements. We use Adam optimizer and a learning rate of 1e61e-6 for both upper and lower-level training. We employ a batch size of $4$ and train on a single rtx6000ada GPU.
For both ICL and MetaICL, during the testing phase the model is given $k=4$ demonstration examples for each test data point. We don’t use demonstration examples in test samples for bilevel-ZOFO evaluation. We evaluate the zero-shot capabilities of our method as well as the performance of the final model LoRA-tuned for 10 additional iterations on $4$ demonstration samples from each class of each test sub-task. Similar to [34], we report Macro-averaged F1 as the evaluation metric.
@@ D.7 Other PEFT Variants
Bilevel-ZOFO is a framework: the lower level can adopt any FO-PEFT method and the upper level any ZO estimator. To illustrate compatibility beyond LoRA-style adapters, we add results with DePT [45], a prompt-tuning method in Table 15. We see gains persist when swapping in a stronger PEFT variant.

### v3

@@ A.2 Fine-tuning LLMs for Multitask and Few-Shot Learning
Multi-task learning (MTL) enables a model to handle multiple tasks simultaneously, fostering knowledge transfer between tasks and improving overall efficiency [MinLZH22MetaICL, yang2024]. Typical meta-tuning approaches employ First-Order methods to train autoregressive LLMs on a multitask dataset for various tasks [ZhongLZK21, MinLZH22MetaICL, GuoXR24, liu2025modality, liu2025explore]. ZhongLZK21 apply meta-training to tasks such as hate speech detection, question categorization, topic classification, and sentiment classification. GuoXR24 adopt the method from MinLZH22MetaICL for generating stylistic text. While MinLZH22MetaICL focus on enhancing the in-context learning ability of the meta-trained model for multitask learning, ZhongLZK21 focus on improving zero-shot performance. This approach is particularly valuable in low-resource settings, where collecting large labeled datasets can be costly, as is often the case with medical data. In such environments, few-shot learning—where a model is fine-tuned on a high-resource dataset to quickly adapt to new tasks with minimal data—becomes essential [YeLR21CrossfitFewShot]. To address the challenges of multi-task and few-shot learning in natural language processing, several meta fine-tuning methods have been proposed [HuLLPDL23, ZhGWZYW24, YeLR21CrossfitFewShot, AsBKLZC24]. However, traditional meta fine-tuning approaches, such as MetaICL [MinLZH22MetaICL], still require full-model first-order gradient calculations, which become computationally expensive with large language models (LLMs) containing billions of parameters. During training, MinLZH22MetaICL sample a task from the dataset for each iteration to perform in-context learning. In contrast to ZhongLZK21 and MinLZH22MetaICL, our approach uses a bilevel structure: the full LLM is fine-tuned at the upper level, while parameter-efficient fine-tuning (PEFT) models are tuned at the lower level. At test time, we freeze the meta-tuned base model and fine-tune only the PEFT model using a few-shot setup, which is both more cost-effective and efficient. Crucially, MinLZH22MetaICL fine tune the full model with first order methods, while we employ a ZO method in meta-tuning the base model at the upper level. Our approach allows us to bypass the need for backpropagation in the meta-model, significantly reducing computational costs.
@@ A.3 First-order Methods for Bilevel Optimization
Solving a bilevel optimization problem is challenging because the function value in the upper-level objective depends on the optimizer of the lower-level problem. This makes it difficult to compute the gradient of the upper-level objective, also known as the hypergradient. Classical methods require calculating Hessian-vector multiplications to approximate the hypergradient [FrDoFrPo17, FrFrSaGrPo18, FiAbLe17, LiGuHu21, RaFiKaLe19, GhWa18, ChKaZh22, LiXiFeZhYoPiUZ18]. However, when fine-tuning large language models, this process becomes extremely expensive due to the high computational and memory demands.
Recently, new frameworks for bilevel optimization have been introduced [LuMei24, ShenC23, LiuLYZZ24, KwonKWN23, LiuYWSL22, ji-bilevel-rebuttal-1, mackay-bilevel-rebuttal-2]. These methods bypass the need for second-order information by reformulating the bilevel problem as a constrained optimization problem. The constraint is penalized, allowing the problem to be tackled as a minimax problem using only first-order information. These methods significantly reduce computational costs by eliminating the need for second-order information. Nevertheless, when fine tuning LLMs, back propagation for calculating the gradient of an LLM is still too expensive.
LiuLYZZ24 and LuMei24 explore the convergence of their proposed methods to the original bilevel problem, while other approaches only demonstrate convergence to the penalized problem. In this paper, we adapt the method from LuMei24 to approximate part of the upper-level parameters using a ZO approximation in order to address the challenge posed by the large number of training parameters in large language models. We also provide convergence guarantees for this adapted zeroth-order-first-order method.
@@ Appendix B Theoretical guarantees of Bilevel ZOFO
The following is the low effective rank assumption from [MalladiGNDL0A23Mezo]. This assumption avoids dimension $d$ in the total complexity. Following [MalladiGNDL0A23Mezo], we assume here that ${\mathbf{z}}^{k}$ in (7) is sampled from shpere in $\mathbb{R}^{d}$ with radius $\sqrt{d}$ for ease of illustration.
@@
This result matches the complexity in previous ZO minimax algorithm in [wang2023zeroth] but solves our bilevel optimization problem (2) and does not depend on the dimensionality $d$ thanks to the efficient rank assumption B.5, providing efficiency guarantee for our algorithm.
@@ C.1 Proofs
(Lemma 1.2.3, Theorem 2.1.8 and Theorem 2.1.10 in [nesterov2013introductory])
@@
The following is a rephrase of theorem 2 in [LuMei24].
@@
Therefore, by Theorem 2 in [LuMei24] we have $\mathbb{E}[\|\nabla F({\bm{\theta}},{\mathbf{p}}^{*}({\bm{\theta}}))\|]\leq O(\epsilon)$ and Lemma B.4 is proven. ∎
@@
(Lemma C.2 in [zhang2024dpzero]) Let $\mathbf{u}$} be uniformly sampled from the Euclidean sphere $\sqrt{d}\mathbb{{\mathbf{s}}}^{d-1}$ and $\mathbf{v}$} be uniformly sampled from the Euclidean ball $\sqrt{d}\mathbb{B}^{d}=\{{\mathbf{x}}\in\mathbb{R}^{d}\mid\|{\mathbf{x}}\|\leq\sqrt{d}\}$. For any function $f({\mathbf{x}}):\mathbb{R}^{d}\to\mathbb{R}$ and $\alpha>0$, we define its zeroth-order gradient estimator as:
@@
The following lemma is from Theorem 1 in [MalladiGNDL0A23Mezo].
@@ D.1 Single-Task experiments
Following MeZO [MalladiGNDL0A23Mezo], we evaluate our approach on a range of classification and multiple-choice tasks: BoolQ [clark2019boolq], CB [wang2019superglue-cb], CB [wang2019superglue-cb], COPA [roemmele2011copa], ReCoRD: [zhang2018record],RTE [wang2018gluesst2], SST2 [wang2018gluesst2], WiC [pilehvar2018wic], WinoGrande [sakaguchi2021winogrande]. In this setting, training and testing are conducted on the same task.
Hyperparameter Search Given resource limitations, we focus on sweeping only the learning rate as the key hyperparameter. For MeZO and first-order PEFT experiments, we explore learning rates from the set $\{1e-2,1e-3,1e-4,1e-5,1e-6\}$. For Bilevel-ZOFO, we sweep both the upper-level and lower-level learning rates: $\text{lr}_{\text{upper}}\in\{1e-4,1e-5,1e-6\}$ and $\text{lr}_{\text{lower}}\in\{1e-2,1e-3,1e-4,1e-5\}$. We perform all experiments in tables 7 and 8 using three random seeds and report the average and standard deviation. We also set $\epsilon=1e-3$, following MeZO [MalladiGNDL0A23Mezo].
@@ D.1.1 Training
All experiments use the Adam optimizer [AdamKingmaB14],including baselines and both lower-level and upper-level optimizers. No weight decay was applied, and the models were trained with a constant learning rate schedule. Batch size is set to $16$ for all experiments. We load all models in bfloat16. We find the best performing model based on validation loss and report test results from that checkpoint. We report the test accuracy or F1-score based on the test dataset being imbalanced or not.
@@ D.2 More Results for single task fine tuning
Table 7 presents the detailed test metrics when applying bilevel-ZOFO and baselines to fine-tune OPT-1.3B [OPT] on a downstream task.
@@
Table 8 demonstrates the results for fine-tuning Llama2-7b [Llama2] on various classification and open-ended generation tasks.
@@ D.2.1 Compare with the two-stage pipeline
The observed performance drop after the second stage is indeed counter-intuitive at first glance. However, it is a limitation of MeZO as it approximates gradients. While further fine-tuning intuitively should improve performance, the inherent noise in gradient approximation can lead to suboptimal updates. This observation is consistent with the fact that MeZO typically requires a significant number of iterations to converge. This is a key contribution of our work: Our approach addresses MeZO’s challenges, such as sensitivity to hard prompts and long convergence times, while outperforming both MeZO and PEFT and maintaining similar memory efficiency. The intuition behind why our method is effective in enhancing both MeZO’s full-model tuning and PEFT is in the nested bilevel structure. This structure encodes more information (as reflected in the training method) from the prompt tuning stage than only treating it as a first stage, thereby providing better guidance for MeZO. In contrast, our bilevel method effectively addresses the issues of MeZO and demonstrates improved performance over both MeZO and the PEFT baseline, even with the same number of ZO iterations. This phenomenon that a bilevel-method is better than a two-staged pipeline is also observed in the later work on diffusion models [shirkavand2024efficient].
The training loss curves for both stages of a two-stage approach and our bilevel framework are provided in Figure 6. When running MeZO in the second stage, the training loss exhibits oscillations and does not show improvement within 500–1000 iterations. This behavior is consistent with findings in the original MeZO [MalladiGNDL0A23Mezo] paper, which notes that MeZO typically requires much longer to converge—on the order of 100k iterations. The oscillatory behavior observed within the shorter training duration is not surprising due to gradient approximation errors.
@@ D.3.1 Meta-Tasks
Following the methodology of MinLZH22MetaICL, we evaluate the performance of bilevel-ZOFO as a fast and efficient meta-learning algorithm. We perform experiments using four of the distinct meta-learning settings outlined in MetaICL [MinLZH22MetaICL]: classification-to-classification, non-classification-to-classification, QA-to-QA, and non-QA-to-QA. Each of these meta-learning tasks includes a set of training sub-tasks and a different set of test sub-tasks. The sub-tasks are sourced from CROSSFIT [YeLR21CrossfitFewShot] and UNIFIEDQA [KhashabiMKSTCH20], comprising a total of 142 unique sub-tasks. These sub-tasks cover a variety of problems, including text classification, question answering, and natural language understanding, all in English. Table 10 shows the number of tasks in each training and testing meta-learning setting and the total number of examples in each training task.
@@
See Tables 14 and 15 of MetaICL [MinLZH22MetaICL] for a list of all sub-tasks.
@@ D.3.2 Baselines
We use GPT2-Large [radford2019gpt2] as the base model for these experiments.We compare our method against several baseline approaches:
  • MetaICL [MinLZH22MetaICL]: A method for meta-learning with in-context learning. MetaICL tunes all the parameters of the base model using the first-order method. In both training and testing, the model is given $k$ demonstration examples, $(a_{1},b_{1}),\dots,(a_{k},b_{k})$}, where $b_{i}$ represents either classification labels or possible answers in question-answering tasks, along with one test example $(a,b)$. The input is formed by concatenating the demonstration examples $a_{1},b_{1},\dots,a_{k},b_{k},a$. The model then computes the conditional probability of each label, and the label with the highest probability is selected as the prediction.
@@
We sample 768 examples from each training sub-task. We use these samples to train MetaICL in their original setting for 30,000 steps. This includes learning rate of 1e51e-5, batch size of $1$ on $8$ GPUs, 8-bit Adam optimizer and fp16 half precision. See MetaICL[MinLZH22MetaICL] for full details. To train our method, we split the training dataset of each sub-task to two subsets, 256 samples as the development dataset for upper-level updates and 512 samples for lower-level training. For each outer iteration of our method, we randomly sample a subset of 5 training tasks. We perform 10 lower-level updates between each pair of upper-level updates. To keep bilevel-ZOFO as lightweight as possible, unlike MetaICL, we do not include demonstration examples in the inputs. Since bilevel-ZOFO uses significantly less memory and has much faster updates compared to MetaICL, theoretically we are able to train it for many more iterations within the same total training duration as MetaICL. However, due to resource constraints, we only train bilevel-ZOFO for 50,000 iterations. Similar to  [MalladiGNDL0A23Mezo], we did not observe a plateau in performance for bilevel-ZOFO, indicating that further training can yield additional improvements. We use Adam optimizer and a learning rate of 1e61e-6 for both upper and lower-level training. We employ a batch size of $4$ and train on a single rtx6000ada GPU.
For both ICL and MetaICL, during the testing phase the model is given $k=4$ demonstration examples for each test data point. We don’t use demonstration examples in test samples for bilevel-ZOFO evaluation. We evaluate the zero-shot capabilities of our method as well as the performance of the final model LoRA-tuned for 10 additional iterations on $4$ demonstration samples from each class of each test sub-task. Similar to [MinLZH22MetaICL], we report Macro-averaged F1 as the evaluation metric.
@@ D.7 Other PEFT Variants
Bilevel-ZOFO is a framework: the lower level can adopt any FO-PEFT method and the upper level any ZO estimator. To illustrate compatibility beyond LoRA-style adapters, we add results with DePT [shi2024dept], a prompt-tuning method in Table 15. We see gains persist when swapping in a stronger PEFT variant.